# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

### v1

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable potential in the field of natural language processing. Unfortunately, LLMs face significant security and ethical risks. Although techniques such as safety alignment are developed for defense, prior researches reveal the possibility of bypassing such defenses through well-designed jailbreak attacks. In this paper, we propose QueryAttack, a novel framework to systematically examine the generalizability of safety alignment. By treating LLMs as knowledge databases, we translate malicious queries in natural language into code-style structured query to bypass the safety alignment mechanisms of LLMs. We conduct extensive experiments on mainstream LLMs, ant the results show that QueryAttack achieves high attack success rates (ASRs) across LLMs with different developers and capabilities. We also evaluate QueryAttack’s performance against common defenses, confirming that it is difficult to mitigate with general defensive techniques. To defend against QueryAttack, we tailor a defense method which can reduce ASR by up to 64% on GPT-4-1106. The code of QueryAttack can be found on https://anonymous.4open.science/r/QueryAttack-334B.
@@ 1 Introduction
Large language models (LLMs) such as OpenAI’s GPT series (OpenAI, 2024a) and Meta’s Llama series (Touvron et al., 2023a) demonstrate remarkable generative potential across various domains (Xiao et al., 2024; Boiko et al., 2023; He et al., 2024; Gao et al., 2024a). However, the immense amounts of data used for training LLMs contain massive information, enabling them to learn unscreened knowledge, including those may evidently violate ethical and moral standards (Li et al., 2023; Jiang et al., 2024; Yuan et al., 2024; Bai et al., 2024). To some extent, LLMs can function as knowledge databases for users. Consequently, a critical responsibility of service providers is to prevent these models from supplying harmful information to potentially adversaries.
@@
Figure 1: The figure illustrates a well-established phenomenon: malicious queries, when translated into a custom language, can bypass the security defenses of LLMs, leading them to generate harmful outputs based on that language. Adversaries can obtain the desired harmful text in natural language following translation. In contrast to these attacks, QueryAttack is capable of directly inducing LLMs to generate harmful text in natural language without requiring a de-translation step.
@@
To align their responses with human ethics and preferences, numerous techniques are employed during the training process of LLMs to regulate their responses to human queries. For example, supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022), reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024), red teaming (Bai et al., 2022a), and the constitutional AI (Bai et al., 2022b) approach are proposed to enhance the safety of LLMs. These efforts enhance the ability of LLMs to identify and defend against malicious queries. Unfortunately, these countermeasures cannot completely prevent outputs that violate human ethics. A significant limitation of these methods is their reliance on malicious language samples from the alignment stage to train the model to recognize malicious queries and ensure the generation of safe outputs. This dependency leaves room for adversaries to develop jailbreak methods using non-natural language as input. For example, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher and Morse code to translate harmful queries into encrypted text. ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding. CodeAttack (Ren et al., 2024b) leverages data structures to embed natural language queries within string variables of programming languages.
We observe that, the essence of these jailbreak attacks lies in defining a customized non-natural language syntax and then using this language to interact with the target LLMs, thereby bypassing their defense mechanisms. To enable the model to understand the customized non-natural language, some methods employ few-shots learning to allow the model to learn how to interact using the specified syntax in context (Peng et al., 2024; Yuan et al., 2024; Barak, 2023; Wei et al., 2023a).
Inspired by prior work, we posit that defenses that relying solely on RLHF cannot achieve a complete defense against jailbreak attacks using non-natural language as inputs. Instead, these defenses can only be incrementally improved by continuously providing LLMs with samples of zero-day attacks. However, this approach is costly and demands sustained effort and resources.
To systematically explore this perspective, we design an effective jailbreak attack to illustrate the challenges of safety generalization in environments that is distant from safety training distribution. Our proposed attack, referred to as QueryAttack, is carried out in three main components: (1) defining a query format based on non-natural language and using a translator to convert natural language queries into the target non-natural language format; (2) prompting the target LLM in context to understand the non-natural language and provide responses in natural language; (3) using the translated query to induce the model to generate harmful responses.
@@
These three steps define a query task, analogous to quering data from a database using code. The additional cost introduced by this process is limited to translating the malicious query into the specified format, which can be easily adapted to any query based on natural languages. Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query format.
We design a total of nine query formats written in the styles of commonly used programming languages and evaluate our attack on both well-known open-source and closed-source LLMs. Experimental results demonstrate that QueryAttack effectively bypasses their security defenses.
@@
  1. 1.
    We propose a hypothesis regarding the safety mechanisms of LLM for natural language alignment, suggesting that solely relying on RLHF cannot prevent the emergence of zero-day attacks, and use it to guide the design of a novel jailbreak attack.
  2. 2.
    We introduce QueryAttack, a novel jailbreak framework for LLMs based on query language translation. QueryAttack only requires the translation of harmful quries without requiring a de-translation step.
  3. 3.
@@
    Our evaluation on mainstream LLMs demonstrates that QueryAttack successfully bypasses their security mechanisms and achieves state-of-the-art average attack success rate (ASR).
2 Related Work
2.1 Jailbreak Attacks on LLMs
@@
Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information. Since natural language samples are widely used as training data during the training phase of LLMs, potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
Initially, researchers reveal that adversaries could launch attacks by manually constructing out-of-distribution (OOD) samples (Li et al., 2023; Shen et al., 2024). However, this kind of approaches relies heavily on extensive human effort, incurs significant costs on launching attacks, or it is difficult to adapt to different models or models trained on different datasets. Building on these observations, several attack methods based on white-box assumptions are proposed (Zou et al., 2023; Jones et al., 2023a, b; Gao et al., 2024c). These attacks leverage access to the model’s architecture and weights of parameters to generate adversarial prompts based on gradients. While effective, these methods are likely to fail when applied to closed-source models under black-box scenarios.
Compared to white-box attacks, black-box attacks have gradually garnered significant attention due to their minimal requirement for model-specific knowledge. In general, black-box attacks assume that adversaries can only adjust their prompt strategies based on the model’s responses (Wei et al., 2023a). Typical black-box attack methods, such as scenario setting-up or role-playing (Zhang et al., 2024; Shah et al., 2023; Liu et al., 2024; Ren et al., 2024c; Song et al., 2025), work by prompting the LLMs into dangerous hypothetical contexts, thereby triggering conflicts with the value alignment mechanisms of these models. Another approach aims to use pre-trained models to adversarially generate harmful prompts by leveraging feedback from the target models as feedback (Chao et al., 2023a).
Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities. By replacing sensitive contents with Base64 (Wei et al., 2023a), ciphertext (Yuan et al., 2024), or low-resource languages (Deng et al., 2024), such attacks induce mismatched generalization in the target LLMs. Notably, some recent work has focused on crafting adversarial inputs using programming languages. For instance,  (Ren et al., 2024b) embed malicious queries within data structures (e.g., stacks and queues) to bypass safety alignments designed for prompts written in natural languages.  (Lv et al., 2024) encrypt malicious prompts using custom functions, transforming them into code completion tasks. The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format. However, these methods typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content.
2.2 Safety Alignment for Defending Jailbreak
To defend against jailbreak attacks, researchers have developed various mitigation strategies to guide large language models in rejecting malicious prompts of potential adversaries.
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is one of the most widely used defense mechanisms. For instance, recent works such as (Mehrabi et al., 2024; Sun et al., 2023) explore the effectiveness of alignment during pre-training in defending against malicious queries, as well as in-context learning (Wei et al., 2023b; Ren et al., 2024a). A common feature of these approaches is the use of red teaming and reinforcement learning (Bai et al., 2022a; Mehrabi et al., 2024). These methods often rely on natural language inputs collected from red teams, which can lead to generalization issues when faced with non-natural language or other OOD inputs.
The latest CoT reasoning-enhanced models, such as OpenAI’s O1 model (OpenAI, 2024c), employ an embedded chain-of-thought mechanism to enforce step-by-step reasoning about the intent of queries and reflection on their compliance with policies during the response generation process. This approach does not rely on text in natural language from the training phase and simultaneously improves the model’s understanding of both queries and policy guidelines.
@@ 3 Methodology
3.1 Observation
As discussed in Section 1, we aim to demonstrate that the current generalization capabilities of safety aligned LLMs are insufficient to defend against zero-day attack samples constructed by adversaries using custom OOD languages. To promote understanding of target models, many jailbreak attack methods employ few-shots or many-shots to help the model learn the user-defined language (Yuan et al., 2024; Jiang et al., 2024). However, custom languages can be highly diverse. The effectiveness of such attacks depends significantly on the model’s ability to understand the new language.
In light of this, some attacks use common but non-natural languages as templates to encrypt malicious queries, achieving high ASR even under zero-shot conditions (Peng et al., 2024; Ren et al., 2024b). These languages often exist within the training samples of LLMs and are well understood by the models. However, they are not fully explored during the safety alignment process, making it possible for the models to generate malicious responses based on harmful prompts written by these languages. Programming languages clearly align with the common characteristics of such non-natural languages. Therefore, in this paper, we use them as examples to construct our query template.
Given the excellent generative capabilities of LLMs, we leverage them to construct translators that converts text in natural language into the our custom query templates. Few-shot learning is employed to guide the model to understand the query intents defined by the templates and generate responses in natural language. For those models with stronger comprehension capabilities (e.g., GPT-O1 (OpenAI, 2024c)), we employ zero-shot learning. Figure 2 provides an overview of QueryAttack.
Figure 2: The overview of QureyAttack. QueryAttack executes a three-step process to jailbreak the target LLM: 1). Translating the harmful query into a code-style custom template. 2. Using in-context learning to help the target LLM understand the natural semantics of the template. 3. Launching the attack using the translated template.
3.2 Query Template
Previous encryption-based jailbreak attacks convert malicious prompts into formats not encountered during the safety alignment phase. CipherChat (Yuan et al., 2024) translates queries into encrypted text using techniques such as Caesar cipher and Morse code. CodeAttack (Ren et al., 2024b) encapsulates entire natural language queries within string variables of programming languages using data structures. CodeChameleon (Lv et al., 2024) designs four different encryption functions to alter the word order of original queries or reconstruct the queries for translation.
These methods require the LLMs to generate responses in a specified non-natural language, which the adversary then decrypts to responses in natural language format. However, these approaches typically depend on the model’s knowledge of the target language to interpret the prompts or place high requirements on the model’s ability to generate text in the target language.
Unlike these attacks, we aim to only translate the malicious prompts into specific formats, but making the target LLMs directly generate responses in natural language. A total of nine commonly used programming languages are selected to design the query templates. Specifically, we identify three query components from natural language queries:
@@ 3.2 Query Template
Using these three components, we define the query templates, as shown in Figure 2. This insight is inspired by the semantics of structured query languages (SQL), which query content from databases using standardized syntax. Similarly, the format of the Hypertext Transfer Protocol Secure (HTTPS) protocol, which uses Uniform Resource Locators (URLs) to fetch web resources from servers, also informs this process. Once the three query components are identified, this query template can be adapted to other programming languages. We provide a total of nine query templates based on commonly used programming languages, which can be found in the appendix A.
3.3 Query Translator
After defining the templates, the primary task is to translate natural language prompts into the template format. Extracting the three query components from the text in natural language is the most critical step in this process. This task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009). Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction.
@@ 3.3 Query Translator
We design specialized prompts to guide the GPT-4-1106 in translation. These prompts are aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model. The details of the prompt are provided in the appendix C.
3.4 Query Learning
@@
The purpose of the query learning is to guide the model in understanding the intent behind the query templates and then generate a natural language response. To help the LLMs understand the content queried by the templates, we first establish the context of the conversation. By describing the three query components, we guide the model in building a mapping from the query template to the natural language, and define the conversation within an educational context.
@@
Few-shot learning is then used to reinforce the model’s understanding of the query and guide it on how to respond to these prompts using natural language. Some text in natural languages, which contain longer modifiers or multiple queries, may require several templates to help define the query. Therefore, we provide both short and long examples.
@@
For models with strong understanding of programming languages, we can skip this process and use zero-shot to launch the attack. Finally, we provide detailed guidance in natural language to respond to the query queries. The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts. In the appendix B, we provide the complete prompt of this part in SQL style version.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable potential in the field of natural language processing. Unfortunately, LLMs face significant security and ethical risks. Although techniques such as safety alignment are developed for defense, prior researches reveal the possibility of bypassing such defenses through well-designed jailbreak attacks. In this paper, we propose QueryAttack, a novel framework to examine the generalizability of safety alignment. By treating LLMs as knowledge databases, we translate malicious queries in natural language into structured non-natural query language to bypass the safety alignment mechanisms of LLMs. We conduct extensive experiments on mainstream LLMs, and the results show that QueryAttack not only can achieve high attack success rates (ASRs), but also can jailbreak various defense methods. Furthermore, we tailor a defense method against QueryAttack, which can reduce ASR by up to 64% on GPT-4-1106. Our code is available at https://github.com/horizonsinzqs/QueryAttack.
@@ 1 Introduction
Large language models (LLMs) such as OpenAI’s GPT series (OpenAI, 2024a) and Meta’s Llama series (Touvron et al., 2023a) demonstrate remarkable generative potential across various domains (Xiao et al., 2024; Boiko et al., 2023; He et al., 2024; Gao et al., 2024a). However, the immense amounts of data used for training LLMs contain massive information, enabling them to learn unscreened knowledge, including those may evidently violate ethical and moral standards (Li et al., 2023; Jiang et al., 2024; Yuan et al., 2024; Bai et al., 2024). Therefore, a critical responsibility of service providers is to prevent these models from supplying harmful information to potentially adversaries.
@@
Figure 1: Existing methods reveal a well-established phenomenon: malicious queries, when well encrypted, can bypass the security defenses of LLMs, leading them to generate encrypted harmful outputs. Unlike these methods, QueryAttack achieves jailbreak by translating malicious inputs into structured non-natural query languages and is capable of directly inducing LLMs to generate harmful content in natural language without the need for a de-translation step.
@@
To align the responses of LLMs with human ethics and preferences, numerous techniques are employed during the training process of LLMs to regulate their outputs to human queries. For example, supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022), reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024), red teaming (Bai et al., 2022a), and the constitutional AI (Bai et al., 2022b) approach are proposed to enhance the safety of LLMs. Unfortunately, a significant limitation of these methods is their reliance on malicious natural language samples from the alignment stage to train the model to recognize malicious queries and ensure the generation of safe outputs. This dependency leaves room for adversaries to develop jailbreak methods using non-natural language as input.
Specifically, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher to translate harmful queries into encrypted text. ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding.  Deng et al. (2024) convert sensitive contents into low-resource languages. The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format. However, they typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content. As a result, their attack effectiveness is limited. Therefore, developing an effective and efficient jailbreak attack method remains a critical challenge.
We observe that, the essence of these jailbreak attacks lies in defining a customized encryption method and then using the language encrypted by this method to interact with the target LLMs, thereby bypassing their defense mechanisms. Inspired by prior work, we find that LLM’s defensive mechanisms are not sensitive to structured, non-natural query languages. For example, by treating the target LLM as a knowledge database, when using structured query language (SQL) to request malicious knowledge (as shown in Figure 1) the target LLM not only identifies the intent of the request well but also does not trigger the defense mechanisms. Instead, the target LLM responds to the entire prompt in natural language normally.
From this new perspective, we propose an attack that first uses structured non-natural query languages to jailbreak LLMs, named QueryAttack. Specifically, we break down QueryAttack into three main components:
1). Extracting three key components from the original query: the requested content, the modifier of the content, and the high-level category to which the content belongs (potential sources where the content can be found).
2). Filling the query components into predefined query templates (e.g., SQL templates) to generate a structured non-natural query.
3). Applying in-context learning to help the target LLM understand the natural semantics of the template and prompting the target LLM using the structured non-natural query.
@@
These three steps define a query task, analogous to querying data from a database using SQL. The additional cost introduced by this process is limited to translating the malicious query into the specified format, which can be easily adapted to any query based on natural languages. Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query templates.
We test QueryAttack on AdvBench (Zou et al., 2023) across both well-known open-source and closed-source LLMs. The experimental results show that QueryAttack effectively bypasses their security defenses, achieving state-of-the-art attack success rates (ASRs). Besides, we provide a visual analysis of QueryAttack’s success and propose a tailored defense method against QueryAttack.
@@
  • We are the first to observe that the defense mechanisms of LLMs are not sensitive to structured non-natural query languages and propose QueryAttack, a novel jailbreak framework based on this observation.
@@
  • Our evaluation on mainstream LLMs demonstrates that QueryAttack successfully bypasses their security mechanisms and achieves state-of-the-art average attack success rate.
  • We propose a tailored defense approach to mitigate QueryAttack, and experiments show that it can effectively help LLMs reduce the attack success rate.
Figure 2: The overview of QureyAttack. QueryAttack executes a three-step process to jailbreak the target LLM: 1). Extracting three key query components from the original query. 2). Filling the query template to get a query code. 3). Applying in-context learning to help the target LLM understand the natural semantics of the template and prompting the target LLM using the query code.
2 Background
@@
Large language models (LLMs) have demonstrated remarkable generative potential across various fields. However, they are still vulnerable to jailbreak attacks. Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information. Since natural language samples are widely used as safety alignment data during the training phase of LLMs (gpt4), potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities. For example, by replacing sensitive contents with Base64 (Wei et al., 2023a), ciphertext (Yuan et al., 2024), or low-resource languages (Deng et al., 2024), such attacks induce mismatched generalization in the target LLMs.
Despite the development of numerous defense methods by researchers to mitigate jailbreaking attacks, such as supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022) and reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024). However, recent work has shown that they are not able to completely defend against zero-day jailbreaking attacks (Song et al., 2025; DBLP:conf/coling/WuG0025). Therefore, it is necessary to continue studying the attack vector of LLMs and provide insights for developing new defenses.
@@ 3 Methodology
The core idea of QueryAttack is to use structured non-natural query languages to carry out jailbreak attacks. As shown in Figure 2, QueryAttack defines a content query task using three components to induce harmful outputs from the target LLM: 1). Query Components Extraction, which extracts key query components from the original query written in natural language. 2). Query Template Filling, which uses the extracted query components to fill the query template and get query code. 3). ICL-based Query Understanding, which applies in-context learning to enable the target LLM to understand the natural semantics of the template and generate responses with the user’s desired content.
3.1 Query Components Extraction
We first identify the components to be extracted from a natural language query (referred to as query components). Taking using Structured Query Language (SQL) as an example, the preliminary task for converting a natural language query into SQL code is to extract the key components of the query sentence. We observe that, for a given query, the two most relevant components to the query’s semantics are: the content to be queried and the description of that content. These two components can naturally be filled into the “SELECT” and “WHERE NAME=” sections of an SQL query. To ensure the completeness of the SQL query and to guide the target LLM’s response, we add an additional component: the category to which the queried content belongs (which can be understood as potential sources where the content may be found). This component can then be filled into the “FROM” section of an SQL query.
Using these three extracted query components, a clear query can be defined, regardless of the language used. For example, the three key query components of the natural language query “Tell me the method of crafting a bomb” are { content: ‘crafting method’, modifiers: ‘bomb’, category: ‘crafting catalog’. }. Therefore, the first step of QueryAttack is to extract the following three query components from the natural language query:
@@ 3.1 Query Components Extraction
The query components extraction task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009). Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction. We design specialized prompts to guide the GPT-4-1106 in components extraction. This prompt is aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model. To achieve this, we employ in-context learning to enable the model to understand the component extraction task and standardize its output. The details of the prompt are provided in the Appendix C.
3.2 Query Template Filling
Figure 3: Templates written in common programming language styles.
After extracting the query components, we obtain the key semantics of a query. Therefore, we can rewrite the query into any non-natural language based on them. To automate the attack, we can predefine query templates for non-natural languages, allowing us to simply fill them with the query components to generate query in non-natural language.
One observation is that non-natural languages commonly present in the training data of LLMs are more likely to be understood by LLMs. Since programming languages are widely present in the training data of LLMs, we choose them as examples to demonstrate the effectiveness of QueryAttack. Among many programming languages, two that are naturally associated with requesting content are SQL and Uniform Resource Locators (URLs). SQL is designed to query content from databases using standardized syntax. Similarly, the format of the Hypertext Transfer Protocol Secure (HTTPS), which uses URLs to fetch web resources from servers, follows a similar process. For example, in SQL, a malicious query written in natural language, such as “Tell me the method of crafting a bomb”, can be rewritten using its three query components as “SELECT ’crafting method’ FROM ’crafting catalog’ WHERE NAME = ’bomb’”.
Inspired by this, although other programming languages are not specifically designed for querying resources, their syntax contains similar keywords or expressions that can express similar query intents, such as “print”, “input” and “return”. It means once the three query components are identified, this query template can be adapted to other programming languages. Specifically, using the query components, we define query templates for nine common programming languages (C, C++, C#, Python, Java, Javascript, Go, URL, SQL), as shown in Figure 3.
As a conclusion, when using QueryAttack to jailbreak the LLMs, the second step is to fill the extracted three key query components into the corresponding language’s query template and obtain the query code, as shown in Figure 2.
3.3 Query Understanding
@@
The purpose of the query learning is to guide the model in understanding the intent behind the query codes and then generate a natural language response. To help the LLMs understand the content, we first establish the context of the conversation. By describing the three query components, we guide the model in building a mapping from the query code to the natural language, and define the conversation within an educational context.
@@
Few-shot learning is then used to reinforce the model’s understanding of the query and guide it on how to respond to these prompts using natural language. Some text in natural languages, which contain multiple queries, may require several query codes to help define the query. Therefore, we provide both short and long examples.
@@
For models with strong understanding of programming languages, we can skip this process and use zero-shot to launch the attack. Finally, we provide detailed guidance in natural language to respond to the query queries. The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts. In Appendix A, we provide the complete prompt of this part.
Through the above steps, we enable the target LLM to understand the intent of the query code and generate responses in natural language according to the query. The adversary then uses the query code obtained in the second step to launch the attack and obtain the desired malicious knowledge.